OpenStudy (anonymous):

HELP ILL MEDAL Juan and Nina are in the Junior High Orchestra. As part of the orchestra they each have to learn two instruments. Juan is learning to play the violin and the cello. For every 3 hours that Juan spends practicing the violin he practices the cello for 2.5 hours. Nina is learning to play the viola and cello. For every 3 hours Nina spends practicing the viola she practices the cello for 2 hours. If Juan practiced the violin for 9 hours last week and Juan and Nina practiced the cello for the same amount of time, then Nina practiced the viola for ________ hours last week.

OpenStudy (anonymous):

Gotta find out what goes in the blank

OpenStudy (anonymous):

J=3/2.5 N=3/2 J practices violin for 9 hours J practices cello for 22.5

OpenStudy (anonymous):

what I have to do to find the answer now ?

OpenStudy (anonymous):

@MGLMystiicz

OpenStudy (anonymous):

I'm so so sorry I have to go but...hey ill be back on later :D

OpenStudy (anonymous):

nvmd , my time gonna expire :(